A coffee merchant has coffee beans that sell for $9 per pound and $12 per pound. The two types are to be mixed to create 100 lb of a mixture that will sell for $11.25 per pound. How much of each of type of bean should be used in the mixture.
Answer by checkley71(8403)

900X+1200(100-X)=1125*100 OR 900X+120000-1200X=112500 OR -300X=7500 OR X=25
OR 25 POUNDS OF 9$ COFFEE WITH 75 POUNDS OF $12 COFFEE MAKES 100 POUINDS OF
$11.25 PER POUND COFFEE
PROOF 900*25+1200(100-25)=1125*100 OR 22500+90000=112500 OR 112500=112500